# Negative cosmological constant in the dark energy sector: tests from JWST photometric and spectroscopic observations of high-redshift galaxies

### v1

@@ 4 Results
We begin by comparing our theoretical predictions against the stellar mass density inferred from the most massive objects identified by NIRCam observations of the JWST CEERS program, as reported in Labbé et al. (2023). In particular, these observations refer to the redshift range $z_{\min}=9\lesssim z\lesssim z_{\max}=11$, which we treat at an effective redshift $z_{\text{eff}}=10$, focusing on the most massive bin considered in Labbé et al. (2023), corresponding to $M_{\star}\geq\overline{M_{\star}}=10^{10.5}M_{\odot}$. As discussed in Sec. 3, from the MCMC chains we extract cosmological parameter vectors consistent within $2\sigma$ with the standard CMB, BAO, and SNeIa cosmological probes, predict the corresponding maximal comoving cumulative stellar mass density, which we then correct for the assumed fiducial cosmology, and then compare this against the JWST CEERS observations. For each model, we derive the exclusion probability $P$, or conversely the probability of consistency with the JWST observations, $Q=1-P$.
@@
The results of our analysis are reported in terms of consistency probability contour plots as a function of the evolving DE parameters $w_{0}$ and $w_{a}$ in Fig. 1, where each of the four quadrants refers to different regimes for the vacuum energy density parameter $\Omega_{\Lambda}$: dS vacuum $\Omega_{\Lambda}>0$ (lower right quadrant), AdS vacuum with $-1<\Omega_{\Lambda}<0$ (lower left quadrant), AdS vacuum with even more negative $\Omega_{\Lambda}<-1$ (upper right panel), and finally any value of $\Omega_{\Lambda}$ (upper left quadrant). Superimposed on the same plots are $2\sigma$ and $3\sigma$ joint confidence intervals in the $w_{0}$-$w_{a}$ plane obtained from a combination of CMB, BAO, and SNeIa measurements (black contours, as reported in Brout et al., 2022), as well as from the combined Hubble diagram of SNeIa and quasars (QSOs, blue contours, as reported in Risaliti & Lusso, 2019), reaching up to redshift $z\sim 5.5$.
@@
The first noteworthy result is that the existence of a region of $w_{0}$-$w_{a}$ parameter space featuring a DE component with an AdS vacuum which is consistent with all the standard cosmological probes and the abundance of high-redshift galaxies inferred from JWST. This can be contrasted to the earlier results of Menci et al. (2022) which only considered a dS vacuum, finding that a major portion of the $w_{0}$-$w_{a}$ parameter space favored by the standard cosmological probes is excluded at significance $>2\sigma$ by the JWST observations. This result on its own, therefore, represents a solid quantitative confirmation of the earlier exploratory results of Adil et al. (2023), and confirms that the presence of a nCC in the DE sector can help accelerate structure formation, thereby aiding the formation of very massive objects at very high redshift beyond what is possible within $\Lambda$CDM (for more detailed explanations of why this occurs, we refer the reader to Adil et al., 2023).
@@
It is also interesting to note that the favored region in $w_{0}$-$w_{a}$ parameter space is not far from ($w_{0}$, $w_{a}$)=($-1$, 00), while lying slightly within the phantom regime, in qualitative agreement with Adil et al. (2023). Within this region, we are able to achieve a consistency probability with the JWST measurements of up to $47\%$. Moreover, this same region is also consistent within $\approx 2\sigma$ with the region favored by SNeIa+QSOs. As $\Omega_{\Lambda}$ is increased, and we therefore move from the upper right quadrant to the two lower quadrants, we see a gradual shift in the required properties of the evolving DE component: this progressively shifts from being characterized by mostly quintessence-like behaviour (potentially crossing the phantom divide in the past, in agreement with the results of Adil et al., 2023), to a quintessence-like behaviour at present but a phantom behaviour in the past, i.e. moving towards the lower right part of the $w_{0}$-$w_{a}$ plane. The latter point is in excellent agreement with the earlier findings of Menci et al. (2022), which indeed focused precisely on this regime, and showed that therein most of the $w_{a}>0$ region is excluded, unless $w_{0}$ is deep in the phantom regime. Either way, it is clear that as $\Omega_{\Lambda}$ is increased, the point ($w_{0}$, $w_{a}$)=($-1$, 00) is progressively disfavored, with a probability of consistency of $<10\%$ in the $\Omega_{\Lambda}>0$ case (lower right quadrant), in agreement with Menci et al. (2022). In the latter case, one should pay attention to the fact that Fig. 2 in Menci et al. (2022) plots the exclusion probability, whereas here we are plotting the probability for consistency: when one accounts for this, it is easy to see that the results are consistent with those of Menci et al. (2022). On the other hand, an AdS vacuum plus evolving DE component scenario displays a good consistency with the JWST measurements (see especially the upper right quadrant): the reason is that the presence of the nCC helps accelerate structure formation, thereby reducing the need for phantom DE, which also happens to be theoretically much harder to accommodate with respect to a quintessence-like component.
@@
While the above conclusions are robust on the theoretical side, critical issues may affect the measurements of Labbé et al. (2023) we are comparing against. In first place, potential uncertainties may affect the calibration of the JWST photometric data used therein. A second issue may concern the Chabrier IMF adopted by Labbé et al. (2023) to derive stellar masses. While we do not expect that assuming other universal forms for the IMF based on low-redshift conditions would change (or even strengthen) the constraints we have derived, the star formation process can be significantly different at redshifts as high as the ones we are concerned with (see e.g. Steinhardt et al., 2023b).
@@
Figure 1: Contours in the $w_{0}$-$w_{a}$ plane colored by the probability of consistency with the stellar mass density inferred from the six most massive galaxies identified in the first NIRCam observations of the JWST CEERS program, in the redshift range $9\lesssim z\lesssim 11$, as reported in Labbé et al. (2023). The contours are reported considering the dynamical DE model discussed in Sec. 2, featuring an evolving DE component with positive energy density, sitting on top of a vacuum energy component with density parameter $\Omega_{\Lambda}\gtrless 0$. The four different quadrants correspond to different regimes for the value of $\Omega_{\Lambda}$: all values $\Omega_{\Lambda}\gtrless 0$ (upper left quadrant), $\Omega_{\Lambda}<-1$ (upper right quadrant), $-1<\Omega_{\Lambda}<0$ (lower left quadrant), and $\Omega_{\Lambda}>0$ (lower right quadrant). As discussed in Sec. 3, we only consider combinations of cosmological parameters consistent within $2\sigma$ with the standard CMB+BAO+SNeIa cosmological probes. Our contours are compared to the $2\sigma$ and $3\sigma$ contours allowed by the latest CMB+BAO+SNeIa measurements (black contours) as reported in Brout et al. (2022), and the combined Hubble diagram of SNeIa and quasars (QSOs, blue contours) reported in Risaliti & Lusso (2019). The green dot corresponds to the $\Lambda$CDM case ($w_{0}=-1$, $w_{a}=0$).
@@
For such reasons, it is important to perform an independent test against spectroscopic data in a different redshift range. With this in mind, we then proceed to carry out a comparison against the stellar mass density inferred from the three most massive, optically dark (dust-obscured) galaxies with robust spectroscopic redshifts from the JWST FRESCO survey (Xiao et al., 2023). Note that here as well the stellar masses have been conservatively inferred using a Chabrier IMF. We expect that assuming other universal forms for the IMF (e.g. Salpeter, Kennicut) would yield similar or even larger values of $M_{\star}$, and therefore our results can be interpreted as being conservative in this sense. As in our previous comparison, we conservatively consider for the measured value $M_{\star}$ the one corresponding to the lower tip of the uncertainty provided by Xiao et al. (2023). As for the uncertainties in the number density, the confidence levels thereon are derived using the statistical method presented in Ebeling (2004), as discussed in Menci et al. (2022).
@@
The results of such a comparison are shown in Fig. 2, with the same quadrant structure as in Fig. 1. We see that the results are completely consistent, and very similar, to those obtained comparing against the photometric sample at higher redshift of Labbé et al. (2023). The only quantitative difference is that overall we are able to achieve higher levels of consistency with the JWST FRESCO data, as high as $98\%$. Moreover, just as observed earlier, we see that the fraction of evolving DE models with phantom behaviour increases with increasing values of $\Omega_{\Lambda}$, confirming the trend seen with the photometric sample: in other words, introducing a nCC helps accelerate structure formation and reduce the need for a phantom DE component, a welcome reduction from the theory point of view.
@@
Figure 2: Same as in Fig. 1, but considering the three most massive, optically dark (dust-obscured) galaxies with robust spectroscopic redshifts identified within the JWST FRESCO survey, in the redshift range $5\lesssim z\lesssim 6$, as reported in Xiao et al. (2023).
@@
For completeness, in Fig. 3 we display the associated contours in the $h$-$S_{8}\equiv\sigma_{8}\sqrt{\Omega_{m}/0.3}$ parameter space, where $h\equiv H_{0}/(100\,{\text{km}}/{\text{s}}/{\text{Mpc}})$ is the reduced Hubble constant, and $S_{8}$ controls the overall strength of matter clustering in the late Universe. While the contours we obtain are, by construction, consistent with the standard cosmological observables we considered, they display some tension with weak lensing (and, to some extent, galaxy clustering) measurements, which favour lower values of $S_{8}$ (see e.g. Di Valentino & Bridle, 2018; Di Valentino et al., 2021d; Nunes & Vagnozzi, 2021). This is of course not unexpected.
@@ 5 Conclusions
The puzzling abundance of extremely massive galaxies at very high redshift unveiled by the early JWST observations has the potential to upturn the current concordance $\Lambda$CDM model, itself already plagued by other observational tensions. At the same time, the JWST observations can be used to test alternative cosmological models, and can potentially rule out those models which do not allow for a sufficiently fast growth of structure required to explain the formation of the galaxies observed by JWST. As argued by various earlier works (Menci et al., 2020, 2022; Adil et al., 2023), observations of the abundance of high-redshift galaxies can place strong constraints on dynamical DE models, in a way which is highly complementary to standard CMB, BAO, and SNeIa cosmological probes at low redshifts. In the earlier exploratory work of Adil et al. (2023) some of us argued that a DE sector featuring a negative cosmological constant (AdS vacuum) with an evolving DE component on top, which carries strong motivation from string theory given the ubiquity of AdS vacua therein, can lead to a more efficient growth of structure at early times, potentially explaining the JWST observations. Our goal in this work has been that of going beyond this exploratory analysis, by a) performing a more careful comparison of such a model against JWST data, while also b) better assessing the complementarity with standard cosmological probes, and c) considering also spectroscopic data from the JWST FRESCO survey (Xiao et al., 2023) in order to place the results on a more solid footing.
Our results, which qualitatively confirm the exploratory findings of Adil et al. (2023), can be summarized as follows:
  • a DE sector featuring an evolving component with positive energy density on top of a negative cosmological constant with density parameter $\Omega_{\Lambda}<0$ can indeed improve consistency with photometric JWST observations at redshifts $9\lesssim z\lesssim 11$ (in agreement with the findings of Adil et al., 2023), with consistency of up to $47\%$;
@@
  • for $\Omega_{\Lambda}>0$, the JWST observations favor dynamical DE models which cross the phantom divide, in agreement with the findings of Menci et al. (2022);
@@
In short, we have shown that DE sectors featuring a negative cosmological constant are extremely interesting from the perspective of the high-redshift galaxies observed by JWST. Moreover, this class of models, and more generally models featuring negative energy densities in the DE sector, also have the potential to partially (albeit not completely) alleviate the Hubble tension (see e.g. Visinelli et al., 2019; Akarsu et al., 2020; Di Valentino et al., 2021a; Acquaviva et al., 2021; Akarsu et al., 2021, 2023a; Sen et al., 2022; Akarsu et al., 2023b; Adil et al., 2023; Akarsu et al., 2023). This therefore confirms that such models are interesting from all three the observational, phenomenological, and also theoretical perspective, in the latter case in light of their strong motivation from string theory.
Overall, our results further strengthen the case for illuminating the nature of DE, and more generally testing new fundamental physics, using observations of high-redshift galaxies, in a redshift range which cannot be reached by standard cosmological probes. We believe such observations have the potential to become an important “emergent probe” (see e.g. Moresco et al., 2022) in coming years, especially with the achievable higher spatial resolution and improved sensitivity from future ALMA/NOEMA and deep JWST spectroscopic observations, whose complementarity with upcoming CMB measurements (e.g. Ade et al., 2019) it would be interesting to explore. From the theory side, it could be interesting to extend our analysis to one implementing a fundamental string scenario from first principles, along the lines of Cicoli et al. (2019) and  Ruchika et al. (2023) (see also Oikonomou et al. (2023)). The early JWST observations have the potential to shake modern cosmology and shed new light onto the nature of the dark sector components. Our work represents a small step in the latter direction, and we cannot wait to see what lies in store for cosmology as our long-awaited space telescope keeps gathering data.
Acknowledgements

### v2

@@ 4 Results
We begin by comparing our theoretical predictions against the stellar mass density inferred from the most massive objects identified by NIRCam observations of the JWST CEERS program, as reported in Ref. [28]. In particular, these observations refer to the redshift range $z_{\min}=9\lesssim z\lesssim z_{\max}=11$, which we treat at an effective redshift $z_{\text{eff}}=10$, focusing on the most massive bin considered in Ref. [28], corresponding to $M_{\star}\geq\overline{M_{\star}}=10^{10.5}M_{\odot}$. As discussed in Sec. 3, from the MCMC chains we extract cosmological parameter vectors consistent within $2\sigma$ with the standard CMB, BAO, and SNeIa cosmological probes, predict the corresponding maximal comoving cumulative stellar mass density, which we then correct for the assumed fiducial cosmology, and then compare this against the JWST CEERS observations. For each model, we derive the exclusion probability $P$, or conversely the probability of consistency with the JWST observations, $Q=1-P$.
@@
The results of our analysis are reported in terms of consistency probability contour plots as a function of the evolving DE parameters $w_{0}$ and $w_{a}$ in Fig. 1, where each of the four quadrants refers to different regimes for the vacuum energy density parameter $\Omega_{\Lambda}$: dS vacuum $\Omega_{\Lambda}>0$ (lower right quadrant), AdS vacuum with $-1<\Omega_{\Lambda}<0$ (lower left quadrant), AdS vacuum with even more negative $\Omega_{\Lambda}<-1$ (upper right panel), and finally any value of $\Omega_{\Lambda}$ (upper left quadrant). Superimposed on the same plots are $2\sigma$ and $3\sigma$ joint confidence intervals in the $w_{0}$-$w_{a}$ plane obtained from a combination of CMB, BAO, and SNeIa measurements (black contours, as reported in Ref. [281]), as well as from the combined Hubble diagram of SNeIa and quasars (QSOs, blue contours, as reported in Ref. [282]), reaching up to redshift $z\sim 5.5$.
@@
The first noteworthy result is that the existence of a region of $w_{0}$-$w_{a}$ parameter space featuring a DE component with an AdS vacuum which is consistent with all the standard cosmological probes and the abundance of high-redshift galaxies inferred from JWST. This can be contrasted to the earlier results of Ref. [94] which only considered a dS vacuum, finding that a major portion of the $w_{0}$-$w_{a}$ parameter space favored by the standard cosmological probes is excluded at significance $>2\sigma$ by the JWST observations. This result on its own, therefore, represents a solid quantitative confirmation of the earlier exploratory results of Ref. [209], and confirms that the presence of a nCC in the DE sector can help accelerate structure formation, thereby aiding the formation of very massive objects at very high redshift beyond what is possible within $\Lambda$CDM (for more detailed explanations of why this occurs, we refer the reader to Ref. [209]).
@@
It is also interesting to note that the favored region in $w_{0}$-$w_{a}$ parameter space is not far from ($w_{0}$, $w_{a}$)=($-1$, 00), while lying slightly within the phantom regime, in qualitative agreement with Ref. [209]. Within this region, we are able to achieve a consistency probability with the JWST measurements of up to $47\%$. Moreover, this same region is also consistent within $\approx 2\sigma$ with the region favored by SNeIa+QSOs. As $\Omega_{\Lambda}$ is increased, and we therefore move from the upper right quadrant to the two lower quadrants, we see a gradual shift in the required properties of the evolving DE component: this progressively shifts from being characterized by mostly quintessence-like behaviour (potentially crossing the phantom divide in the past, in agreement with the results of Ref. [209]), to a quintessence-like behaviour at present but a phantom behaviour in the past, i.e. moving towards the lower right part of the $w_{0}$-$w_{a}$ plane. The latter point is in excellent agreement with the earlier findings of Ref. [94], which indeed focused precisely on this regime, and showed that therein most of the $w_{a}>0$ region is excluded, unless $w_{0}$ is deep in the phantom regime. Either way, it is clear that as $\Omega_{\Lambda}$ is increased, the point ($w_{0}$, $w_{a}$)=($-1$, 00) is progressively disfavored, with a probability of consistency of $<10\%$ in the $\Omega_{\Lambda}>0$ case (lower right quadrant), in agreement with Ref. [94]. In the latter case, one should pay attention to the fact that Fig. 2 in Ref. [94] plots the exclusion probability, whereas here we are plotting the probability for consistency: when one accounts for this, it is easy to see that the results are consistent with those of Ref. [94]. On the other hand, an AdS vacuum plus evolving DE component scenario displays a good consistency with the JWST measurements (see especially the upper right quadrant): the reason is that the presence of the nCC helps accelerate structure formation, thereby reducing the need for phantom DE, which also happens to be theoretically much harder to accommodate with respect to a quintessence-like component.
@@
While the above conclusions are robust on the theoretical side, critical issues may affect the measurements of Ref. [28] we are comparing against. In first place, potential uncertainties may affect the calibration of the JWST photometric data used therein. A second issue may concern the Chabrier IMF adopted by Ref. [28] to derive stellar masses. While we do not expect that assuming other universal forms for the IMF based on low-redshift conditions would change (or even strengthen) the constraints we have derived, the star formation process can be significantly different at redshifts as high as the ones we are concerned with (see e.g. Ref. [283]).
@@
Figure 1: Contours in the $w_{0}$-$w_{a}$ plane colored by the probability of consistency with the stellar mass density inferred from the six most massive galaxies identified in the first NIRCam observations of the JWST CEERS program, in the redshift range $9\lesssim z\lesssim 11$, as reported in Ref. [28]. The contours are reported considering the dynamical DE model discussed in Sec. 2, featuring an evolving DE component with positive energy density, sitting on top of a vacuum energy component with density parameter $\Omega_{\Lambda}\gtrless 0$. The four different quadrants correspond to different regimes for the value of $\Omega_{\Lambda}$: all values $\Omega_{\Lambda}\gtrless 0$ (upper left quadrant), $\Omega_{\Lambda}<-1$ (upper right quadrant), $-1<\Omega_{\Lambda}<0$ (lower left quadrant), and $\Omega_{\Lambda}>0$ (lower right quadrant). As discussed in Sec. 3, we only consider combinations of cosmological parameters consistent within $2\sigma$ with the standard CMB+BAO+SNeIa cosmological probes. Our contours are compared to the $2\sigma$ and $3\sigma$ contours allowed by the latest CMB+BAO+SNeIa measurements (black contours) as reported in Ref. [281], and the combined Hubble diagram of SNeIa and quasars (QSOs, blue contours) reported in Ref. [282]. The green dot corresponds to the $\Lambda$CDM case ($w_{0}=-1$, $w_{a}=0$).
@@
For such reasons, it is important to perform an independent test against spectroscopic data in a different redshift range. With this in mind, we then proceed to carry out a comparison against the stellar mass density inferred from the three most massive, optically dark (dust-obscured) galaxies with robust spectroscopic redshifts from the JWST FRESCO survey [29]. Note that here as well the stellar masses have been conservatively inferred using a Chabrier IMF. We expect that assuming other universal forms for the IMF (e.g. Salpeter, Kennicut) would yield similar or even larger values of $M_{\star}$, and therefore our results can be interpreted as being conservative in this sense. As in our previous comparison, we conservatively consider for the measured value $M_{\star}$ the one corresponding to the lower tip of the uncertainty provided by Ref. [29]. As for the uncertainties in the number density, the confidence levels thereon are derived using the statistical method presented in Ref. [280], as discussed in Ref. [94].
@@
The results of such a comparison are shown in Fig. 2, with the same quadrant structure as in Fig. 1. We see that the results are completely consistent, and very similar, to those obtained comparing against the photometric sample at higher redshift of Ref. [28]. The only quantitative difference is that overall we are able to achieve higher levels of consistency with the JWST FRESCO data, as high as $98\%$. Moreover, just as observed earlier, we see that the fraction of evolving DE models with phantom behaviour increases with increasing values of $\Omega_{\Lambda}$, confirming the trend seen with the photometric sample: in other words, introducing a nCC helps accelerate structure formation and reduce the need for a phantom DE component, a welcome reduction from the theory point of view.
@@
Figure 2: Same as in Fig. 1, but considering the three most massive, optically dark (dust-obscured) galaxies with robust spectroscopic redshifts identified within the JWST FRESCO survey, in the redshift range $5\lesssim z\lesssim 6$, as reported in Ref. [29].
We stress that the CEERS and FRESCO datasets adopted are not the only JWST-based measurements of the abundance of high-redshift galaxies. The reason why we have adopted these and not others, however, is that at present they constitute the most constraining datasets, and are therefore best suited for the scope of the present work. For instance, NIRCam observations of the GLASS-ERS 1324 program reported in Ref. [284] involve galaxies at comparable redshifts but with much lower stellar masses (log $M_{\star}/M_{\odot}\lesssim 9.5$), and are therefore not expected to significantly tighten existing cosmological constraints. On the other hand, JWST observations of the Hubble Ultra Deep Field (HUDF) and UKIDSS Ultra Deep Survey field reported in Ref. [285] extend to lower redshifts ($z\lesssim 8$) compared to ones we considered. Indeed, the measurements reported in Ref. [285] yield number densities which are comparable to, or even lower than, those reported in Ref. [286] for the CANDELS/UDS, GOODS-South, and HUDF fields. The cosmological implications of the measurements reported in Ref. [286] were analyzed by one of us in Ref. [88], finding much weaker constraints on dynamical DE compared to the ones obtained in Ref. [94] based on JWST CEERS data, and in the present work. Therefore, we do not expect that the measurements presented in Refs. [284, 285] would add much constraining power to constraints on dark energy from standard cosmological probes, although a study thereof could be interesting for follow-up works. The CEERS and FRESCO measurements we consider in the present work provide at present an unprecedented combination of high redshifts, large masses, and large abundances which make them uniquely suited for our study.
@@
For completeness, in Fig. 3 we display the associated contours in the $h$-$S_{8}\equiv\sigma_{8}\sqrt{\Omega_{m}/0.3}$ parameter space, where $h\equiv H_{0}/(100\,{\text{km}}/{\text{s}}/{\text{Mpc}})$ is the reduced Hubble constant, and $S_{8}$ controls the overall strength of matter clustering in the late Universe. While the contours we obtain are, by construction, consistent with the standard cosmological observables we considered, they display some tension with weak lensing (and, to some extent, galaxy clustering) measurements, which favour lower values of $S_{8}$ (see e.g. Refs. [287, 288, 289]). This is of course not unexpected.
@@
Before closing, a comment is in order regarding our treatment of the cosmic baryon fraction $f_{b}$. As explained earlier, we have not treated $f_{b}$ as a free parameter, but as a derived parameter which is self-consistently computed from the values of $\omega_{b}$, $\Omega_{m}$, and $H_{0}$ at each point in the MCMC chains. This approach differs from a number of other studies on the cosmological implications of the JWST measurements, where $f_{b}$ is either fixed to the best-fit value inferred within $\Lambda$CDM from the Planck 2018 measurements, or freely varied within some prior range. We believe our approach is somewhat more robust as it ensures that $f_{b}$ is self-consistently constrained by the other probes at play, varying consistently with the same parameters which control our predictions for the growth of structure and thereby the stellar mass density. Of course, it is worth noting that the inferred values of $f_{b}$ are model-dependent, and can be affected by the choice of both pre- and post-recombination models. Indeed, it is known that the assumed pre-recombination model can affect the inferred value of $\Omega_{m}h^{2}$, a fact which has been appreciated in the context of the Hubble tension [290], whereas $\Omega_{b}h^{2}$ is essentially fixed by considerations on Big Bang Nucleosynthesis and the relative height of the odd versus even acoustic peaks in the CMB. We note that higher (lower) values of $f_{b}$ would lead to proportionally higher (lower) values of the cumulative comoving number and stellar mass densities of galaxies (as both scale linearly with $f_{b}$), thereby slightly weakening (strengthening) our existing constraints. Nevertheless, within the range of allowed model-dependent values of $f_{b}$, the effect is expected to be small, and changes to the growth history associated to a different dark energy model are expected to be much larger.
@@ 5 Conclusions
The puzzling abundance of extremely massive galaxies at very high redshift unveiled by the early JWST observations has the potential to upturn the current concordance $\Lambda$CDM model, itself already plagued by other observational tensions. At the same time, the JWST observations can be used to test alternative cosmological models, and can potentially rule out those models which do not allow for a sufficiently fast growth of structure required to explain the formation of the galaxies observed by JWST. As argued by various earlier works [88, 94, 209], observations of the abundance of high-redshift galaxies can place strong constraints on dynamical DE models, in a way which is highly complementary to standard CMB, BAO, and SNeIa cosmological probes at low redshifts. In the earlier exploratory work of Ref. [209] some of us argued that a DE sector featuring a negative cosmological constant (AdS vacuum) with an evolving DE component on top, which carries strong motivation from string theory given the ubiquity of AdS vacua therein, can lead to a more efficient growth of structure at early times, potentially explaining the JWST observations. Our goal in this work has been that of going beyond this exploratory analysis, by a) performing a more careful comparison of such a model against JWST data, while also b) better assessing the complementarity with standard cosmological probes, and c) considering also spectroscopic data from the JWST FRESCO survey [29] in order to place the results on a more solid footing.
Our results, which qualitatively confirm the exploratory findings of Ref. [209], can be summarized as follows:
  • a DE sector featuring an evolving component with positive energy density on top of a negative cosmological constant with density parameter $\Omega_{\Lambda}<0$ can indeed improve consistency with photometric JWST observations at redshifts $9\lesssim z\lesssim 11$ (in agreement with the findings of Ref. [209]), with consistency of up to $47\%$;
@@
  • for $\Omega_{\Lambda}>0$, the JWST observations favor dynamical DE models which cross the phantom divide, in agreement with the findings of Ref. [94];
@@
In short, we have shown that DE sectors featuring a negative cosmological constant are extremely interesting from the perspective of the high-redshift galaxies observed by JWST. Moreover, this class of models, and more generally models featuring negative energy densities in the DE sector, also have the potential to partially (albeit not completely) alleviate the Hubble tension (see e.g. Refs. [191, 291, 193, 292, 293, 294, 195, 295, 197, 296]). This therefore confirms that such models are interesting from all three the observational, phenomenological, and also theoretical perspective, in the latter case in light of their strong motivation from string theory.
Overall, our results further strengthen the case for illuminating the nature of DE, and more generally testing new fundamental physics, using observations of high-redshift galaxies, in a redshift range which cannot be reached by standard cosmological probes. We believe such observations have the potential to become an important “emergent probe” (see e.g. Refs. [10]) in coming years, especially with the achievable higher spatial resolution and improved sensitivity from future ALMA/NOEMA and deep JWST spectroscopic observations, whose complementarity with upcoming CMB measurements [297, 298] it would be interesting to explore. From the theory side, it could be interesting to extend our analysis to one implementing a fundamental string scenario from first principles, along the lines of Refs. [299, 192, 300]). The early JWST observations have the potential to shake modern cosmology and shed new light onto the nature of the dark sector components. Our work represents a small step in the latter direction, and we cannot wait to see what lies in store for cosmology as our long-awaited space telescope keeps gathering data.
Acknowledgments